&flushed O Not Secure WHAT WILL HAPPEN I A Mamadou Dialo& Spring 2019 ECON 2106-078 (CRN 17654) Homework: Chapter 32 Homework (Micro) conomicScore: 0 of 1 p Hw Score: 7083%, as of 12pts 4 of 12 (11 complete) X Concept Question 1.24 Queetion Help The tollowing table shown production ponsibilites for Melayse and Ecuador, assuming that here is no trade. Each mumberpenls units of pd hour, all resources are fully employed prouced per Suppose that workers in both nations work a ten-hour day and spil their ime evenly between producing the two goods Total world output of shoes will be 13600 Enter your บายwer its eer fees and then doi Check Answer. All parts showing Clear All ths course Spring 2019 ECON 2106-07(CR 176540 s bused on Miller Economics Today 19 20 2
Answer #1

Malaysia produces 500 MP3 players in 1 hour and Ecuador produces 400 MP3 players in an hour. So, total production of MP3 players in five hours (workers produce MP3 players half of the time) = 5*500 + 5*400 = 2,500 + 2000 = 4,500.

Ecuador produces 900 shoes in 1 hour and Ecuador produces 1,800 shoes in an hour. So, total production of shoes in five hours (workers produce shoes half of the time) = 5*900 + 5*1,800 = 4,500 + 9,000 = 13,500.

Malaysia produces 500 MP3 players in 1 hour. Therefore, it will produce 500 * 10 = 5,000 MP3 players when it specializes in MP3 players.

Ecuador produces 1,800 shoes in 1 hour. Therefore, it will produce 1,800 * 10 = 18,000 shoes when it specializes in shoes.